OpenStudy (anonymous):

Solve. –2(6 – 20n) = 20(2n – 3) A. 40 B. –36 C. no solution D. all real numbers

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

because -12+40n=40n-60

OpenStudy (anonymous):

oooooohhhhhh..

OpenStudy (anonymous):

Determine which equation below has a solution of x = 2. A. 5x + 8 – 12x = –6x + 2x – 3 B. 6x + 4 – 12x = –10x + 18x – 3 C. –4x – 6 + x = –2x + x – 10 D. –3x – 8 + x = –4x + x – 10